The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method.

Answer by stanbon(75887)

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method.
--------------------
Equations:
5x + 2y = 18
3x - 7y = -22
----------
Modify for elimination::
15x + 6y = 3*18
15x - 35y= 5*-22
----------------
Subtract and solve for "y"::
41y = 164
y = 4
------
Solve for "x"::
3x - 7y = -22
3x - 28 = -22
3x = 6
x = 2
------------
Cheers,
Stan H.
---------------